A firm faces the following linear inverse demand for its product P = 60 - 2Q. a)Find the firm's total revenue function TR (Q).b)Find the expression for the firm's marginal revenue.c)Assumingthat the marginal cost of production is given by MC=8. What will be the equilibrium output and price





1
Expert's answer
2021-07-21T14:23:21-0400

a)

"Total \space Revenue=Price\times Quantity"

"TR=P\times Q\\

TR=(60-2Q)Q\\

TR=60Q-2Q^2"

The total revenue function of the firm is :"TR(Q)=60Q-2Q^2"


b) Marginal Revenue:It is the additional revenue obtained by selling an extra output.

Marginal Revenue=Change in Total Revenue"\div" Change in Output

For this, we will differentiate the Total Revenue(TR) function with respect to output (Q).

"MR=\frac{dTR}{dQ}\\

MR=\frac{d(60Q-2Q^2)}{dQ}\\

MR=60-4Q"

The expression for the firm's marginal revenue is "MR=60-4Q"

 

c) Given: Marginal Cost(MC)=8

At the equilibrium,

Marginal Revenue=Marginal Cost

"MR=MC\\

60-4Q=8\\

4Q=60-8\\

4Q=52\\

Q=\frac{52}{4}\\

Q=13"

The equilibrium output is 13

 Putting value of equilibrium output in the demand function to get the equilibrium price

"P=60-2Q\\

P=60-2(13)\\

P=60-26\\

P=34"

The equilibrium price is 34